OpenStudy (anonymous):

Help my brain is overloading!!! Given f(x)= x^2-2x-1 find f(x^2+3) I got x^4-2x^2+2 but its x^4-4x^2+2

OpenStudy (anonymous):

can you tell me why its 4 rather than 2?

OpenStudy (anonymous):

probably algebra

OpenStudy (anonymous):

this problem has been bothering me for the last 2 hours im so frustrated

OpenStudy (anonymous):

\[f(x)= x^2-2x-1\]\[ f(x^2+3) =(x^2+3)^2-2(x^2+3)-1\]

OpenStudy (anonymous):

it isnt worth that much of your time

OpenStudy (anonymous):

yea thats exactly what i did and i got my answer

OpenStudy (anonymous):

course i didn't do it, maybe your answer is correct shall we check?

OpenStudy (anonymous):

but i don't understand why its a 4x^2 rather than 2x^2

OpenStudy (anonymous):

\[(x^2+3)^2=x^4+6x^2+9\]

OpenStudy (anonymous):

when you subtract the \(2x^2\) from the second term you will be left with \(4x^2\) did you forget to square correctly?

OpenStudy (anonymous):

why is it 6x^2

OpenStudy (anonymous):

tsk tks, you deserve to lose two hours if you forgot how do square something please don't tell me you think \((a+b)^2=a^2+b^2\) when it is really \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

wait but why do you do 2ab tho? i don't understand .... its not like we're factoring anything here

OpenStudy (anonymous):

you don't "do" anything, you multiply \[(x^2+3)^2=(x^2+3)(x^2+3)=x^4+3x^2+3x^2+9=x^4+6x^2+9\]

OpenStudy (anonymous):

wow......

OpenStudy (anonymous):

okay im super stupid....

OpenStudy (anonymous):

wow...... *facepalm* okay thanks ...

OpenStudy (anonymous):

yw